OpenStudy (anonymous):

If the length of a rectangle is eight more than thrice the width..write a simplified equation for its perimeter.

OpenStudy (agent0smith):



OpenStudy (anonymous):

so w+3w+8 is the equation?

OpenStudy (anonymous):



OpenStudy (anonymous):

that means 2w+2(3w+8) is the exact equation?

OpenStudy (anonymous):

So it would be \(2(3w+8+w)=2(4w+8)=8w+16\)

OpenStudy (anonymous):

It would be the (3w+8+w) you wrote, but since that's only on of the length plus one of the width, you can just multiply that by 2! Which is what I did haha

OpenStudy (anonymous):

Either way, it simplifies to the same!

OpenStudy (anonymous):

hahahaha..thanks a lot:-D

OpenStudy (anonymous):

No prob man